OpenStudy (anonymous):

How do you simplify the ratio 30mL:45mL

OpenStudy (dumbcow):

reduce the fraction 30/45 = 2/3

OpenStudy (callisto):

30mL:45mL = 30mL/45mL = (5*2*3 )/(3*3*5) = 2/3 = 2:3

OpenStudy (s):

Divide both by 15 and get 2mL : 3mL

OpenStudy (anonymous):

Thanks